Six pounds of grass seed covers 150 square feet. My lawn is 200 square feet. How much do I need
Answer by nerdybill(7384)

Let x = amount (pounds) needed
then
6/150 = x/200
cross multiplying
150x = 1200
x = 1200/150
x = 8 pounds